OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - y = -11 2x + 3y = 5

OpenStudy (zale101):

Have you solved systems of linear equations before? If so, do you have any ideas of where to start?

OpenStudy (anonymous):

no i have idea

OpenStudy (anonymous):

no idea**

OpenStudy (zale101):

4x - y = -11 2x + 3y = 5 Here's the steps to these kind of questions. Step 1: Choose one of the equation above and solve for one variable x or y (depends on your choice) Step 2: After you solve for, let's say X, you plug whatever you got for x into the second equation that you didn't began with. Step 3: When you plug in what you solved for x into the second equation, all the variables will become y. You solve for y, and you'll get a number. Plug that number to either one of the equations and see what you'll get for x. If none of that seems familiar with you, then i will do my best to help and make you understand this.

OpenStudy (zale101):

Step 1: Which equation are you choosing from: ? 4x - y = -11 2x + 3y = 5

OpenStudy (anonymous):

does it matter which one i choose

OpenStudy (zale101):

Nope. It does not matter.

OpenStudy (zale101):

Which equation is the easiest to solve for y?

OpenStudy (zale101):

I'd go with the easiest and i'll choose the first equation.

OpenStudy (anonymous):

2x+3y=5 there is no negatives so its probably easier to solve

OpenStudy (zale101):

Okay, we'll go with that.

OpenStudy (anonymous):

oops lol we can go with yours

OpenStudy (zale101):

Okay.

OpenStudy (zale101):

i'll choose the variable y because it's alone by itself. We can get rid of negatives when we divide both sides by negative :)

OpenStudy (zale101):

\(\large 4x - y = -11 \) Subtract both sides by 4x \(\large 4x-4x - y = -11-4x \) \(\large - y = -11-4x \) Divide both sides by -1 \(\Large \frac{- y}{-1} = \frac{-11-4x}{-1} \) \(\large y = 11+4x \)

OpenStudy (zale101):

Makes sense?

OpenStudy (zale101):

what is y equal to now?

OpenStudy (anonymous):

14x ? im sorry im so confused on this, and my teacher isnt responding

OpenStudy (zale101):

Why 14 x? We can't add 11 with 4x because they are different terms. Y is equal to 11+4x, that's the only info we know

OpenStudy (anonymous):

idk what to do

OpenStudy (zale101):

What does step 2 say?

OpenStudy (anonymous):

After you solve for, let's say X, you plug whatever you got for x into the second equation that you didn't began with.

OpenStudy (zale101):

Yes. But let's not deal with x since we are dealing with y. So, in different wording, we take whatever we got for y and plug it for the y in the second equation. Makes sense?

OpenStudy (zale101):

y=11+4x Our second equation is 2x+3y=5 \(\large 2x+3(11+4x)=5 \)

OpenStudy (anonymous):

2x+33+12x=5 ? or do you not use distributive property

OpenStudy (zale101):

Yes, you already distributed it

Vocaloid (vocaloid):

yup, you're right so far, now let's keep going right now we have: 2x+33+12x=5 now add the x's together: 2x + 12x = __x fill in the blank

OpenStudy (anonymous):

how do i solve that? add or multiply?

OpenStudy (anonymous):

so 14x

OpenStudy (anonymous):

oh yeah sorry

Vocaloid (vocaloid):

yes, so now we have 14x + 33 = 5 subtract 33 from each sign

Vocaloid (vocaloid):

*from each side

OpenStudy (anonymous):

-19=-28

OpenStudy (anonymous):

am i supposed to subtract 33 from 33 too or no

Vocaloid (vocaloid):

not quite, remember that you can't combine variables and numbers 14x + 33 - 33 = 14x 5 - 33 = -28 14x = -28, make sense so far?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

14x = -28 divide both sides by 14

OpenStudy (anonymous):

x=-2

Vocaloid (vocaloid):

good, we're almost done we know that x = -2 now, so we can find y let's go back to our first equation: 4x - y = -11 can you tell me what y is now?

OpenStudy (anonymous):

-2

Vocaloid (vocaloid):

not quite. let's try plugging in -2 into the equation 4*(-2) - y = -11 -8 - y = -11 now can you tell me what y is?

OpenStudy (anonymous):

no i dont understand

Vocaloid (vocaloid):

-8 - y = -11 add 8 to both sides, what do you get?

OpenStudy (anonymous):

-y=-19

Vocaloid (vocaloid):

add, not subtract -11 + 8 = ?

OpenStudy (anonymous):

-3

Vocaloid (vocaloid):

right, so -y = -3 what is y?

OpenStudy (anonymous):

3

Vocaloid (vocaloid):

boom, we're done x = -2 and y = 3 our solution is (-2,3)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -2 can you help with this one too?